Required information [The following information applies to the questions displayed below.] Laker Company reported the following January purchases and sales data for its only product. Date Activities Units Acquired at Cost Units sold at Retail Jan. 1 Beginning inventory 195 units @ $ 12.00 = $ 2,340 Jan. 10 Sales 155 units @ $ 21.00 Jan. 20 Purchase 120 units @ $ 11.00 = 1,320 Jan. 25 Sales 135 units @ $ 21.00 Jan. 30 Purchase 290 units @ $ 10.50 = 3,045 Totals 605 units $ 6,705 290 units The Company uses a perpetual inventory system. For specific identification, ending inventory consists of 315 units, where 290 are from the January 30 purchase, 5 are from the January 20 purchase, and 20 are from beginning inventory. Required: 1. Complete the table to determine the cost assigned to ending inventory and cost of goods sold using specific identification. 2. Determine the cost assigned to ending inventory and to cost of goods sold using weighted average. 3. Determine the cost assigned to ending inventory and to cost of goods sold using FIFO. 4. Determine the cost assigned to endinginventory and to cost of goods sold using LIFO.
Answer #1

Preparing Table of Quantity for opening, purchase, Sales & Closing Stock

Date Opening Purchase COGS Closing
Unit Unit Unit Unit
Jan.1 195 195
Jan.10 195 155 40
Jan.20 40 120 160
Jan.25 160 135 25
Jan.30 25 290 315

1. Determination of the cost assigned to ending inventory & cost of goods sold using specific identification -

Ending inventory cost = 290 qty. are from January 30 purchase (290*10.5) + 5 qty. are from january20 purchase (5*11) + 20 qty.are from beginning inventory (20*12)

= 290*10.5 + 5*11 + 20*12

= 3045 + 55 + 240

= 3340

COGS = opening + total purchases - Ending Inventory

= 2340 + 4365 - 3340 ( as per above calculation)

COGS = 6705 - 3340

= 3365

2. Determination of the cost assigned to ending inventory & cost of goods sold using Weighted Average -

Date Opening Purchase COGS Closing
Unit Rate Value Unit Rate Value Unit Rate Value Unit Rate Value
Jan.1 195 12 2340 195 12 2340
Jan.10 195 12 2340 155 12 1860 40 12 480
Jan.20 40 12 480 120 11 1320 160 11.25 1800
Jan.25 160 11.25 1800 135 11.25 1518.75 25 11.25 281.25
Jan.30 25 11.25 281.25 290 10.5 3045 315 10.56 3326.25

Ending Inventory as per above table 315 units @ 10.56 = 3326.25

COGS = 155*12 + 135*11.25

= 1860 + 1518.75

= 3378.75

3. Determination of the cost assigned to ending inventory & cost of goods sold using FIFO -

Date Opening Purchase COGS Closing
Unit Rate Value Unit Rate Value Unit Rate Value Unit Rate Value
Jan.1 195 12 2340 195 12 2340
Jan.10 195 12 2340 155 12 1860 40 12 480
Jan.20 40 12 480 40 12 480
Jan.20 120 11 1320 120 11 1320
Jan.25 40 12 480 40 12 480 0 0 0
Jan.25 120 11 1320 95 11 1045 25 11 275
Jan.30 25 11 275 25 11.00 275
Jan.30 290 10.5 3045 290 10.5 3045

Ending Inventory Cost = 25 units @ 11 + 290 units @ 10.5

= 25*11 + 290*10.5

= 275 + 3045

= 3320

COGS = 155 units @ $ 12 + 40 units @ $ 12 + 95 units @ $ 11

= 1860 + 480+ 1045

= 3385

4. Determination of the cost assigned to ending inventory & cost of goods sold using LIFO -

Date Opening Purchase COGS Closing
Unit Rate Value Unit Rate Value Unit Rate Value Unit Rate Value
Jan.1 195 12 2340 195 12 2340
Jan.10 195 12 2340 155 12 1860 40 12 480
Jan.20 40 12 480 40 12 480
Jan.20 120 11 1320 120 11 1320
Jan.25 40 12 480 15 12 180 25 12 300
Jan.25 120 11 1320 120 11 1320 0 0 0
Jan.30 25 12 300 25 12.00 300
Jan.30 290 10.5 3045 290 10.5 3045

Ending Inventory Cost = 25 units @ 12 + 290 units @ 10.5

= 25*12 + 290*10.5

= 300 + 3045

= 3345

COGS = 155 units @ $ 12 + 15 units @ $ 12 + 120 units @ $ 11

= 1860 + 180+ 1320

= 3360

Please check with your answer and let me know.